A 6.3 gram sample of an isotope of strotium- 90 was formed in a 1962 explosion of an atomic bomb. The half- life of strontium- 90 is 28 years. In what year will only .07 grams remain?
Answer by scott8148(6628)

.07 = 6.3 (1 / 2)^((x - 1962) / 28)

log(.07 / 6.3) = ((x - 1962) / 28)[log(1 / 2)]

{[log(.07 / 6.3)] / [log(1 / 2)]} * 28 = x - 1962

[{[log(.07 / 6.3)] / [log(1 / 2)]} * 28] + 1962 = x